A radioactive isotope has an activity of 9.25×104 Bq initially. After 3.15 hours the activity is 6.94×104 Bq. What is the half-life of the isotope?

What is the activity after an additional 3.15 hours?
Answer #1

A) let the half life is T

NOw, for the activity

A = Ao * 0.50^(t/T)

6.94 *10^4 = 9.25 *10^4 * 0.50^(3.15/T)

T = 7.6 hours

the half life is 7.6 hours

B)

NOw, for the activity

after 3.15 hours

A = Ao * 0.50^(t/T)

A = 6.94 *10^4 * 0.50^(3.15/7.6)

A = 5.2 *10^4 decays/s

the activity after an additional is 5.2 *104 decays/s